# SAMBA-ASR State-of-the-Art Speech Recognition Leveraging Structured State-Space Models

### v1

@@ Abstract
We propose Samba-ASR, the first state-of-the-art Automatic Speech Recognition (ASR) model leveraging the novel Mamba architecture as both encoder and decoder, built on the foundation of state-space models (SSMs). Unlike transformer-based ASR models, which rely on self-attention mechanisms to capture dependencies, Samba-ASR effectively models both local and global temporal dependencies using efficient state-space dynamics, achieving remarkable performance gains. By addressing the limitations of transformers, such as quadratic scaling with input length and difficulty in handling long-range dependencies, Samba-ASR achieves superior accuracy and efficiency.
Experimental results demonstrate that Samba-ASR surpasses existing open-source transformer-based ASR models across various standard benchmarks, establishing it as the new state-of-the-art in ASR. Extensive evaluations on the benchmark dataset show significant improvements in Word Error Rate (WER), with competitive performance even in low-resource scenarios. Furthermore, the inherent computational efficiency and parameter optimization of the Mamba architecture make Samba-ASR a scalable and robust solution for diverse ASR tasks.
Our contributions include:
  • A new Samba-ASR architecture for ASR, demonstrating SSM’s superiority over transformer-based models for speech sequence processing. .
  • A comprehensive evaluation on public benchmarks showcasing SOTA performance.
  • In-depth analysis of computational efficiency, robustness to noise, and sequence generalization.
This work highlights the viability of Mamba-SSMs as a transformer-free alternative for efficient and accurate ASR. By leveraging the advancements of state-space modeling, Samba-ASR redefines ASR performance standards and sets a new benchmark for future research in this field.

### v2

# Samba-ASR: State-Of-The-Art Speech Recognition Leveraging Structured State-Space Models

@@ Abstract
We propose Samba ASR, the first state-of-the-art Automatic Speech Recognition (ASR) model leveraging the novel Mamba architecture as both encoder and decoder, built on the foundation of state-space models (SSMs). Unlike transformer-based ASR models, which rely on self-attention mechanisms to capture dependencies, Samba ASR effectively models both local and global temporal dependencies using efficient state-space dynamics, achieving remarkable performance gains. By addressing the limitations of transformers, such as quadratic scaling with input length and difficulty in handling long-range dependencies, Samba ASR achieves superior accuracy and efficiency.
Experimental results demonstrate that Samba ASR surpasses existing open-source transformer-based ASR models across various standard benchmarks, establishing it as the new state-of-the-art in ASR. Extensive evaluations on the benchmark dataset show significant improvements in Word Error Rate (WER), with competitive performance even in low-resource scenarios. Furthermore, the inherent computational efficiency and parameter optimization of the Mamba architecture make Samba ASR a scalable and robust solution for diverse ASR tasks.
Our contributions include the development of a new Samba ASR architecture for automatic speech recognition (ASR), demonstrating the superiority of structured state-space models (SSMs) over transformer-based models for speech sequence processing. We provide a comprehensive evaluation on public benchmarks, showcasing state-of-the-art (SOTA) performance, and present an in-depth analysis of computational efficiency, robustness to noise, and sequence generalization. This work highlights the viability of Mamba SSMs as a transformer-free alternative for efficient and accurate ASR. By leveraging the advancements of state-space modeling, Samba ASR redefines ASR performance standards and sets a new benchmark for future research in this field.